OpenStudy (anonymous):

You use a scale of 1 inch = 280 feet to build a model of the Great Pyramid of Giza. The width of the base of the pyramid is 756 feet. What is the width of your model? (1 point)27 in. 3.4 in. 3 in. 2.7 in.

OpenStudy (anonymous):

proportion time ... 1/280 = x/756 (1 inch to 280 ft = x in to 756 ft)-cross multiply (280)(x) = (1)(756) can you take it from here ?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

your welcome :)